How would I calculate the radius of a cricle with a circumference of 3pi? And how would you work it out. Thank you very much
Answer by nerdybill(7384)

How would I calculate the radius of a cricle with a circumference of 3pi?
.
First, we KNOW that the circumference of a circle is:
circumference = (pi)d
where d is the diameter
And, since
r = d/2 (half the diameter)
.
so now, we can say:
circumference = (pi)(2r)
.
the problem gave the circumference as 3(pi) -- plugging it in:
3(pi) = (pi)(2r)
notice the "pi" cancels leaving:
3 = (2r)
3/2 = r (radius)